OpenStudy (chillout):

Show that if a square matrix satisfies A²-3A+I=0, it is nonsingular and its inverse is equal to A^(-1)=3I-A.

OpenStudy (chillout):

LaTex not showing correctly. Editing question.

OpenStudy (vishweshshrimali5):

umm... okay let me give you a hint.. \(A^2 - 3A + I = 0\) \(\implies A^2 - 3A = -I\) \(\implies A(A - 3I) = -I\) \(\implies A^{-1} A (A-3I) = A^{-1} (-I)\)

OpenStudy (vishweshshrimali5):

So can you find \(A^{-1}\) from here ?

OpenStudy (chillout):

Yeah, I was just assuming a 2x2 matrix and doing all the dirty work. You put it plain simple. Thanks!

OpenStudy (vishweshshrimali5):

:) Glad to be of help

OpenStudy (chillout):

My browser is a bit messed up. One moment while I restart it.

OpenStudy (chillout):

\(A^{-1}\cdot A=I \\I(A-3I)=-A^{-1}\rightarrow (A-3I)=-A^{-1} \\\rightarrow A-3I=-A^{-1},\ \ A^{-1}=3I-A\)

OpenStudy (vishweshshrimali5):

Great! Also, for an inverse to exist, the matrix must be non - singular so that proves it

OpenStudy (chillout):

Thanks. Just remembering some linear algebra and I'm so bad with this xD